OpenStudy (anonymous):

the average number of red, blue, and green beanbags in a store is 136. there are 30 more red beanbags than blue beanbags. there are 15 fewer green beanbags then blue beanbags. how many green beanbags are in the store

OpenStudy (anonymous):

fewer green beanbags of what blue or red

OpenStudy (anonymous):

fewer green than blue? sorry edited the question.

OpenStudy (mathstudent55):

Let r = number red beanbags b = number blue beanbags g = number green beanbags From the total number you know that r + b + g = 136 from "there are 30 more red beanbags than blue beanbags" you know r = b + 30 and from "there are 15 fewer green beanbags then blue beanbags." you know g = b - 15 Now substitute what r and g are equal to in the last two equyations into the first equation and solve for b. Then substitute the value of b into the 2nd and 3rd equations to get r and g.

OpenStudy (anonymous):

still having trouble with problem blue equals what?

OpenStudy (mathstudent55):

r + b + g =136 make both substitutions for r and g and you get b + 30 + b + b - 15 = 136 3b + 15 = 336 3b = 321 b = 107